OpenStudy (anonymous):

how do i substitute 6 for x in 210x^2

OpenStudy (campbell_st):

its \[210 \times 6^2 = \]

OpenStudy (mathstudent55):

Where you see x in the expression, replace it with 6. \(210\color{red}{x}^2\) \(=210\times \color{red}{6}^2\)

OpenStudy (anonymous):

Thanks

OpenStudy (mathstudent55):

If you need to evaluate the expression for x = 6, now square the 6 , then multiply by 210.

OpenStudy (mathstudent55):

You're welcome.